OpenStudy (anonymous):

What is the solution to the equation? 3/4(4c+16)=2c+9? c= ?

OpenStudy (anonymous):

What I usually like to do is turn the fraction into what it really is, which is multiplying the equation on the left by 3 and then dividing it by 4. 3*(4c+16) = 12c + 48 Then divide by 4, so we get: 12c/4 + 48/4 3c + 12 So that's on the left side 3c + 12 = 2c + 9 Then let's get the c's on the left 3c - 2c + 12 = 2c - 2c + 9 1c + 12 = 9 Then get the 12 on the other side 1c + 12 - 12 = 9 - 12 1c = -3 c= -3